# Can black holes preserve $\mathcal{N}>4$ supersymmetry?

### v3

@@ 1 Introduction and motivation
[…]
@@
  • $\mathcal{N}=2$ Schwarzian - [$\text{SU}(1,1|1)$], with $\text{U}(1)$ $R$-symmetry and a complex fermion. This theory typically appears for near-BPS black holes in AdS Boruch:2022tno ; Castro:2021wzn ; Maulik:2024dwq ; Heydeman:2024ezi ; Heydeman:2024fgk ; Cabo-Bizet:2024gny and makes predictions about the BPS and near-BPS spectrum of the dual superconformal field theories.
  • “Small” $\mathcal{N}=4$ Schwarzian - $[\text{PSU}(1,1|2)]$, with an $\text{SU}(2)$ $R$-symmetry and a complex fermion in the fundamental representation. As an example, this theory appears for near-BPS black holes in flat space as well as black holes in the $AdS_{3}$ limit of the D1-D5 system Heydeman:2020hhw . This is relevant for justifying Strominger-Vafa style microstate counting Strominger:1996sh by showing the longstanding issue raised in Preskill:1991tb is circumvented by a supersymmetric mass gap.
@@
  • $\mathcal{N}=3$ Schwarzian - $[\text{OSp}(3|2)]$, with $\text{SO}(3)$ $R$-symmetry and a fermion in the vector representation. Depending on global considerations, the quantum $R$-symmetry could be $\text{SU}(2)$, $\text{SO}(3)$ or an anomalous $\text{SO}(3)$.
  • “Large” $\mathcal{N}=4$ Schwarzian - [D$(2,1|\upalpha)$], with $\text{SU}(2)_{+}\times\text{SU}(2)_{-}$ $R$-symmetry and a complex fermion in the bi-fundamental. The lower scripts are simply to distinguish the two groups. This theory additionally comes with a continuous parameter $\upalpha$ and the model depends sensitively on this.
@@
  • $\mathcal{N}=n$ Schwarzian - [$\text{OSp}(n|2)$ for $n\geq 4$], with several choices. At the quantum level there are three versions of the theory, depending on global considerations. They are distinguished by the $R$-symmetry being $\text{SO}(n)$ (anomalous or non-anomalous) and $\text{Spin}(n)$. In the latter case, the fermion fields in the Schwarzian are in vector representations, but this theory has black hole states transforming in spinor representations. [^4] In the 1d Schwarzian or the 2d CFT context, a theory based on $\text{OSp}(2|4)$ is not the $\upalpha\to 1$ limit of D$(1,2|\upalpha)$. This is only true when the $\text{OSp}(2|4)$ theory is quantized such that the $R$-symmetry group is $\text{Spin}(4)$ and not $\text{SO}(4)$. In 2d CFT this appears as a choice in how to sum over integer spectral flow.
  • $\mathcal{N}=7$ Schwarzian - [G$(3)$], with $R$-symmetry group G2 and fermions transforming in the 7-dimensional fundamental of G2.
  • $\mathcal{N}=8$ Schwarzian - [F$(4)$], which has the $R$-symmetry group $\text{Spin}(7)$ (note that F$(4)$ is the superconformal algebra and not the $R$-symmetry). The fermions transforms in the spinor representation of $\text{Spin}(7)$.
@@ 2.2 The action and the partition function
where $\alpha\cdot H_{R}$ is short for $\sum_{i=1}^{\text{rank}\,G_{R}}\alpha_{i}H_{R}^{i}$. We can quantize the theory either using fermionic localization Stanford:2017thb or using the canonical approach of Mertens:2017mtv . Even though the second approach connects more with the work of Knizhnik:1986wc , we follow the localization method which is simpler for our purposes. The integration space for all these theories is a coadjoint orbit of the associated 1d superconformal group, and therefore is a symplectic space. Moreover, we assume that the supersymmetric completion of the Schwarzian action is unique such that it also acts as a generator of a $\text{U}(1)$ symmetry on the supermanifold. This is true for actions generated by $BF$ theories described above. We can then apply the Duistermaat-Heckman theorem and compute the path integral by localization on fixed points of this $\partial/\partial\tau$. In other words,
@@
    Now we turn to the fluctuations around these fixed points. Using the Cartan-Weyl basis for the Lie algebra $\mathfrak{g}_{R}$, the fluctuations can be parametrized by dim$(R)$ functions $\epsilon(\tau)$’s as:
@@
×$\times\exp\left[S_{0}+\frac{2\pi^{2}\Phi_{r}}{\beta}\left(1-8q\,{\rm Tr}_{\rho}\left[\left((\alpha+\nu)\cdot H_{R}\right)^{2}\right]\right)\right].$
@@ 2.3 Some examples
×$\times\frac{e^{{S_{0}+\frac{2\pi^{2}\Phi_{r}}{\beta}(1-\frac{4(1+\upalpha)}{\upalpha}(\alpha_{+}+n)^{2}-4(1+\upalpha)(\alpha_{-}+m)^{2})}}}{(1-4(\alpha_{+}+\alpha_{-}+m+n)^{2})(1-4(\alpha_{+}-\alpha_{-}-m+n)^{2})}.$
@@
×$\quad\times\,e^{\frac{2\pi^{2}\Phi_{r}}{\beta}\left(1-\sum_{i=1}^{l}16(\alpha_{i}+n_{i}/2)^{2}\right)}$
@@ 3.1 Two theories with $\text{SO}(3)$ $R$-symmetry and their solution
[…]
@@ 3.2 Commutation relations of conserved charges
The conserved charges are the Hamiltonian $H$, three supercharges $Q^{i}$ with $i=1,2,3$, transforming in the adjoint of $\text{SO}(3)$ generated by the remaining charges $J^{i}$. We could derive their commutation relations via an explicit construction of the charges from the super Schwarzian action, but instead of following this tedious route, the information we have so far uniquely determines the algebra. The $\text{SO}(3)$ generators satisfy $[J^{i},J^{j}]={\rm i}\varepsilon_{ijk}\,J^{k}$, and since $Q^{i}$ are in the adjoint of $\text{SO}(3)$ they satisfy $[J^{i},Q^{j}]={\rm i}\varepsilon_{ijk}\,Q^{k}$. Both of these generators should commute with the Hamiltonian $[J^{i},H]=[Q^{i},H]=0$. We also define an operator $(-1)^{\sf F}$ such that it commutes with $H$ and $J^{i}$ and anticommutes with $Q^{i}$.
@@
where $\upgamma_{1}$ and $\upgamma_{2}$ are free parameters. We still need to check Jacobi identities. When evaluating [$[\{Q_{i},Q_{j}\},Q_{k}]+(\text{cyclic})$ we find it only vanishes when $\upgamma_{1}=0$. All values of $\upgamma_{2}$ are allowed so we found a one-parameter nonlinear correction of the algebra. The value of that parameter depends on the theory, and in our case it can depend on $\Phi_{r}$. We will see below that an analysis of the BPS spectrum leads to the following identification
[…]
@@
The action of $J^{-}$ generates an $\text{SO}(3)$ multiplet with spin $j$. The action of $Q^{3}$ generates another spin $j$ multiplet of opposite fermion parity. A similar pair of spin $j-1$ multiplets and opposite fermion parity can be constructed by further acting with $Q^{-}$, [^10] To ensure orthogonality within the supermultiplet, the bottom-left state in (37) requires the linear combination of $Q^{-}|E,j\rangle$ and $J^{-}Q^{3}|E,j\rangle$, while the bottom-right state requires linear combination of $Q^{-}Q^{3}|E,j\rangle$ and $J^{-}|E,j\rangle$. See Schoutens:1988tg for more details. making a total multiplet $2(j)\oplus 2(j-1)$. This can be represented in a diagram
@@
This gives us the correct ground state energy with the inverse square root edge found in (30). There is really no gap within the non-BPS multiplets, but there are also no BPS states except for the special case of $j=0$! This inverse edge implies the bound is not satisfied for $j\neq 0$ in gravity, but we can still contemplate what happens when it is. Interestingly, when the BPS bound is saturated, it means that the action of $Q^{3}$ becomes trivial, but not necessarily $Q^{-}$. The diagram is
$j$$j-1$$Q^{-}$$|E_{0},j\rangle$ (39)
@@
$[Q^{+},J^{-}]=Q^{3}.$ (40)
Since $j=0$ we have $J^{-}|E,j=0\rangle=0$ and being a highest-weight state $Q^{+}|E,j=0\rangle=0$. This implies that $[Q^{+},J^{-}]|E,j=0\rangle=0$ and therefore $Q^{3}|E,j=0\rangle=0$. Acting on $\text{SO}(3)$ singlets we have the relation $\{Q^{3},Q^{3}\}=H$ which implies that when $j=0$, $E=0$. Finally, for a state annihilated by $J^{i}$ and $Q^{3}$ the commutation relation $[J^{\pm},Q^{3}]=\pm{\rm i}Q^{\pm}$ implies that $Q^{-}$ also annihilates the state, rendering it fully BPS. Could we have BPS states that transform nontrivially under $\text{SO}(3)$ but also preserve all supercharges? The answer is no since
@@
[…]
@@ 3.3 The index for $\mathcal{N}=3$
The Witten index of the model involves an insertion of $(-1)^{\sf F}$ in the partition function. The list above shows that this insertion not only removes non-BPS states but also removes partially BPS states. Therefore, the Witten index is sensitive only to fully BPS states. We can also use a different operator, $e^{{\rm i}\pi J^{3}}$, which also leads to vanishing contributions from both non- and partially-BPS states. Notice that $(-1)^{\sf F}\neq e^{{\rm i}\pi J^{3}}$ since $[Q^{3},J^{3}]=0$ while $\{Q^{3},(-1)^{\sf F}\}=0$. The two possible indices we can define are
@@ 3.4 Linearization of the model, its solution, and the index
This implies that when $\alpha=1/2$ the partition function is again protected. The non-BPS characters are $02\cdot 2\cdot(\chi_{j}+\chi_{j-1})=$ 0 while the BPS characters are $02\cdot(\chi_{j}+\chi_{j-1})=$ 0 for partially BPS states and $22\cdot\chi_{0}=$ 2 for fully BPS states. Therefore we conclude
@@
Therefore $Z_{\rm linear}(\beta,\alpha=1/2)$ is a protected quantity capturing fully-BPS states. (As far as we know we cannot define a protected quantity capturing partial BPS states.) The explanation is clear, since $[J^{i},\tilde{Q}]=0$ while $\{(-1)^{\sf F},\tilde{Q}\}=0$, this again implies that $(-1)^{\sf F}\neq e^{{\rm i}\pi J}$. In the context of holographic 2d CFT this construction is lifted to the modified elliptic genus introduced in Eberhardt:2018sce .
@@ 4.1 Solution of the model
×$\times\frac{e^{{S_{0}+\frac{2\pi^{2}\Phi_{r}}{\beta}(1-\frac{4}{\gamma_{-}}(\alpha_{+}+n)^{2}-\frac{4}{\gamma_{+}}(\alpha_{-}+m)^{2})}}}{(1-4(\alpha_{+}+\alpha_{-}+m+n)^{2})(1-4(\alpha_{+}-\alpha_{-}-m+n)^{2})}.$
[…]
@@
with the index $A$ and $\dot{A}$ transforming in the bifundamental of $\text{SU}(2)_{+}$ and $\text{SU}(2)_{-}$, respectively. We can pick a basis such that $Q_{1\dot{B}}$ ($Q_{2\dot{B}}$) has eigenvalue $+1/2$ ($-1/2$) with respect to $J_{+}{}^{3}$, and similarly for the second index. We can also describe the $\text{SU}(2)$ generators in the same language as $J_{A}{}^{B}=(\sigma^{i})_{A}{}^{B}J_{+}^{i}$ and $J_{\dot{A}}{}^{\dot{B}}=(\sigma^{i})_{\dot{A}}{}^{\dot{B}}J_{-}^{i}$, where $\sigma^{i}$ are the Pauli matrices.
@@
×$\times\sinh\left(2\pi\sqrt{2\Phi_{r}\left(E-\frac{\upalpha j_{+}^{2}+j_{-}^{2}}{2\Phi_{r}(1+\upalpha)}\right)}\right),$
@@
To describe the allowed spins of the BPS states more explicitly, consider multiplets with fixed total spin $j_{+}+j_{-}=j$ and work with one value of $j$ at a time. For fixed $j$ we have $12j+$ 1 ways of distributing the total spin between the two $\text{SU}(2)$ generators, namely
@@
We find that for a given $j$ there is a unique choice of $(j_{+},j_{-})$ carrying all BPS states, but the precise way the total spin is distributed among the two $\text{SU}(2)$ spins depends on the parameter $\upalpha$. For $\upalpha\sim 0$ our expression for $R_{\rm BPS}$ implies that the only multiplets with BPS states can be $(j_{+}=j,j_{-}=0)$. When we start raising $\upalpha$ there is a jump at $\upalpha=1/(2j)$ since the combination that appears in $R_{\rm BPS}$ approaches its upper bound. At $\upalpha=1/(2j)$ the representation with BPS states jumps from $(j,0)$ to $(j-1/2,1/2)$. As we continue to increase $\upalpha$ we encounter a finite number of jumps at
@@
This happens until we reach $\upalpha=2j$ and we transition from $(1/2,j-1/2)$ to $(0,j)$ at $\upalpha=2j$. For any $\upalpha>2j$ the BPS spin is $(0,j)$. We emphasize that for a fixed value of $\upalpha$ and fixed $j=j_{+}+j_{-}$ there is a unique BPS spin. Finally, we can check that the transition values $\upalpha_{*}$ take place precisely at the zeros of (66).

### v4

@@ 1 Introduction and motivation
There is a loose analogy between the effective field theory of near extremal black holes as dictated by (super)symmetry principles and the possible symmetries of the interacting S-matrix. The Coleman-Mandula theorem Coleman:1967ad famously shows that the possible symmetries of the S-matrix must factorize as $\textbf{Poincare}\times\textbf{Internal}$, which is analogous to the $\text{SL}(2,\mathbb{R})\times G$ symmetry implied by the work of Kunduri, Lucietti and Reall. Of course, this theorem has a famous “loophole” in the form of the Haag–Łopuszański–Sohnius theorem Haag:1974qh , which allows for nontrivial mixing of $\textbf{Poincare}\times\textbf{Internal}$ by the addition of fermionic generators. For near-BPS black holes, a similar mixing of near-horizon symmetries is possible via fermionic isometries, and can be related to the classification of superconformal groups. We will show how this leads to new interesting interacting theories and classify the possibilities.
@@
  • $\mathcal{N}=2$ Schwarzian - $\text{SU}(1,1|1)$], with $\text{U}(1)$ $R$-symmetry and a complex fermion. This theory typically appears for near-BPS black holes in AdS Boruch:2022tno ; Castro:2021wzn ; Maulik:2024dwq ; Heydeman:2024ezi ; Heydeman:2024fgk ; Cabo-Bizet:2024gny and makes predictions about the BPS and near-BPS spectrum of the dual superconformal field theories.
  • “Small” $\mathcal{N}=4$ Schwarzian - $\text{PSU}(1,1|2)]$, with an $\text{SU}(2)$ $R$-symmetry and a complex fermion in the fundamental representation. As an example, this theory appears for near-BPS black holes in flat space as well as black holes in the $AdS_{3}$ limit of the D1-D5 system Heydeman:2020hhw . This is relevant for justifying Strominger-Vafa style microstate counting Strominger:1996sh by showing the longstanding issue raised in Preskill:1991tb is circumvented by a supersymmetric mass gap.
@@
  • $\mathcal{N}=3$ Schwarzian - $\text{OSp}(3|2)]$, with $\text{SO}(3)$ $R$-symmetry and a fermion in the vector representation. Depending on global considerations, the quantum $R$-symmetry could be $\text{SU}(2)$, $\text{SO}(3)$ or an anomalous $\text{SO}(3)$.
[…]
@@
[…]
  • $\mathcal{N}=7$ Schwarzian - [G(3)(3)], with $R$-symmetry group G2 and fermions transforming in the 7-dimensional fundamental of G2.
  • $\mathcal{N}=8$ Schwarzian - [F(4)(4)], which has the $R$-symmetry group $\text{Spin}(7)$ (note that F(4)(4) is the superconformal algebra and not the $R$-symmetry). The fermions transforms in the spinor representation of $\text{Spin}(7)$.
@@ 2.2 The action and the partition function
where $\alpha\cdot H_{R}$ is short for $\sum_{i=1}^{\text{rank}\,G_{R}}\alpha_{i}H_{R}^{i}$. We can quantize the theory either using fermionic localization Stanford:2017thb or using the canonical approach of Mertens:2017mtv . Even though the second approach connects more with the work of Knizhnik:1986wc , we follow the localization method which is simpler for our purposes. The integration space for all these theories is a coadjoint orbit of the associated 1d superconformal group, and therefore is a symplectic space. Moreover, we assume that the supersymmetric completion of the Schwarzian action is unique such that it also acts as a generator of a $\text{U}(1)$ symmetry on the supermanifold. This is true for actions generated by $BF$ theories described above [^7] The BF approach is used simply to systematically generate supersymmetric completions of the Schwarzian action, but otherwise this perspective is not very important for this paper.. We can then apply the Duistermaat-Heckman theorem and compute the path integral by localization on fixed points of this $\partial/\partial\tau$. In other words,
@@
    Now we turn to the fluctuations around these fixed points. Using the Cartan-Weyl basis for the Lie algebra $\mathfrak{g}_{R}$, the fluctuations can be parametrized by dim(R)(R) functions $\epsilon(\tau)$’s as:
@@
$\times\exp\left[S_{0}+\frac{2\pi^{2}\Phi_{r}}{\beta}\left(1-8q\,{\rm Tr}_{\rho}\left[\left((\alpha+\nu)\cdot H_{R}\right)^{2}\right]\right)\right].$
@@ 2.3 Some examples
$\times\frac{e^{{S_{0}+\frac{2\pi^{2}\Phi_{r}}{\beta}(1-\frac{4(1+\upalpha)}{\upalpha}(\alpha_{+}+n)^{2}-4(1+\upalpha)(\alpha_{-}+m)^{2})}}}{(1-4(\alpha_{+}+\alpha_{-}+m+n)^{2})(1-4(\alpha_{+}-\alpha_{-}-m+n)^{2})}.$
@@
$\quad\times\,e^{\frac{2\pi^{2}\Phi_{r}}{\beta}\left(1-\sum_{i=1}^{l}16(\alpha_{i}+n_{i}/2)^{2}\right)}$
@@ 3.1 Two theories with $\text{SO}(3)$ $R$-symmetry and their solution
$Z_{\nu}=e^{S_{0}}\delta_{\nu,0}+\sum_{j\in\mathbb{Z}+\frac{\nu}{2},\,j>0}\int_{E_{0}(j)}^{\infty}{\rm d}E\,e^{-\beta E}\,2(\chi_{j}(\alpha)+\chi_{j-1}(\alpha))\,\rho_{j}(E),$ (28)
@@ 3.2 Commutation relations of conserved charges
The conserved charges are the Hamiltonian $H$, three supercharges $Q^{i}$ with $i=1,2,3$, transforming in the adjoint of $\text{SO}(3)$ generated by the remaining charges $J^{i}$. We could derive their commutation relations via an explicit construction of the charges from the super Schwarzian action, but instead of following this tedious route, the information we have so far uniquely determines the algebra. The $\text{SO}(3)$ generators satisfy $J^{i},J^{j}]={\rm i}\varepsilon_{ijk}\,J^{k}$, and since $Q^{i}$ are in the adjoint of $\text{SO}(3)$ they satisfy $J^{i},Q^{j}]={\rm i}\varepsilon_{ijk}\,Q^{k}$. Both of these generators should commute with the Hamiltonian $J^{i},H]=[Q^{i},H]=0$. We also define an operator $(-1)^{\sf F}$ such that it commutes with $H$ and $J^{i}$ and anticommutes with $Q^{i}$.
@@
where $\upgamma_{1}$ and $\upgamma_{2}$ are free parameters. We still need to check Jacobi identities. When evaluating $\{Q_{i},Q_{j}\},Q_{k}]+(\text{cyclic})$ we find it only vanishes when $\upgamma_{1}=0$. All values of $\upgamma_{2}$ are allowed so we found a one-parameter nonlinear correction of the algebra. The value of that parameter depends on the theory, and in our case it can depend on $\Phi_{r}$. We will see below that an analysis of the BPS spectrum leads to the following identification
$\{Q_{i},Q_{j}\}=\delta_{ij}H-\frac{1}{16\Phi_{r}}\{J_{i},J_{j}\}~~~~\text{(Correct)}.$ (34)
@@
The action of $J^{-}$ generates an $\text{SO}(3)$ multiplet with spin $j$. The action of $Q^{3}$ generates another spin $j$ multiplet of opposite fermion parity. A similar pair of spin j1j-1 multiplets and opposite fermion parity can be constructed by further acting with $Q^{-}$, [^11] To ensure orthogonality within the supermultiplet, the bottom-left state in (37) requires the linear combination of $Q^{-}|E,j\rangle$ and $J^{-}Q^{3}|E,j\rangle$, while the bottom-right state requires linear combination of $Q^{-}Q^{3}|E,j\rangle$ and $J^{-}|E,j\rangle$. See Schoutens:1988tg for more details. making a total multiplet $2(j)\oplus 2(j-1)$. This can be represented in a diagram
[…]
@@
This gives us the correct ground state energy with the inverse square root edge found in (30). There is really no gap within the non-BPS multiplets, but there are also no BPS states except for the special case of j=0j=0! This inverse edge implies the bound is not satisfied for $j\neq 0$ in gravity, but we can still contemplate what happens when it is. Interestingly, when the BPS bound is saturated, it means that the action of $Q^{3}$ becomes trivial, but not necessarily $Q^{-}$. The diagram is
[…]
@@
$Q^{+},J^{-}]=Q^{3}.$ (40)
Since $j=0$ we have $J^{-}|E,j=0\rangle=0$ and being a highest-weight state $Q^{+}|E,j=0\rangle=0$. This implies that $Q^{+},J^{-}]|E,j=0\rangle=0$ and therefore $Q^{3}|E,j=0\rangle=0$. Acting on $\text{SO}(3)$ singlets we have the relation $\{Q^{3},Q^{3}\}=H$ which implies that when $j=0$, $E=0$. Finally, for a state annihilated by $J^{i}$ and $Q^{3}$ the commutation relation $J^{\pm},Q^{3}]=\pm{\rm i}Q^{\pm}$ implies that $Q^{-}$ also annihilates the state, rendering it fully BPS. Could we have BPS states that transform nontrivially under $\text{SO}(3)$ but also preserve all supercharges? The answer is no since
@@
and similarly for $Q^{+}$. If the state preserves all supercharges then the LHS would vanish, but the RHS cannot vanish unless j=0j=0121212Another possibility is that $j=1/2$ since now $J^{-}J^{-}|+1/2\rangle=0$. Only in this representation $\{J^{i},J^{j}\}$ is proportional to $\delta_{ij}$. This case is special since even the non-BPS multiplet only includes two copies of $(1/2)$ while a BPS one (either fully or only preserving $Q^{3}$) only has one copy of $(1/2)$..
@@ 3.3 The index for $\mathcal{N}=3$
The Witten index of the model involves an insertion of $(-1)^{\sf F}$ in the partition function. The list above shows that this insertion not only removes non-BPS states but also removes partially BPS states. Therefore, the Witten index is sensitive only to fully BPS states. We can also use a different operator, $e^{{\rm i}\pi J^{3}}$, which also leads to vanishing contributions from both non- and partially-BPS states. Notice that $(-1)^{\sf F}\neq e^{{\rm i}\pi J^{3}}$ since $Q^{3},J^{3}]=0$ while $\{Q^{3},(-1)^{\sf F}\}=0$. The two possible indices we can define are
@@ 3.4 Linearization of the model, its solution, and the index
This implies that when $\alpha=1/2$ the partition function is again protected. The non-BPS characters are $2\cdot 2\cdot(\chi_{j}+\chi_{j-1})=0$ while the BPS characters are $2\cdot(\chi_{j}+\chi_{j-1})=0$ for partially BPS states and $2\cdot\chi_{0}=2$ for fully BPS states. Therefore we conclude
@@
Therefore $Z_{\rm linear}(\beta,\alpha=1/2)$ is a protected quantity capturing fully-BPS states. (As far as we know we cannot define a protected quantity capturing partial BPS states.) The explanation is clear, since $J^{i},\tilde{Q}]=0$ while $\{(-1)^{\sf F},\tilde{Q}\}=0$, this again implies that $(-1)^{\sf F}\neq e^{{\rm i}\pi J}$. In the context of holographic 2d CFT this construction is lifted to the modified elliptic genus introduced in Eberhardt:2018sce .
@@ 4.1 Solution of the model
$\times\frac{e^{{S_{0}+\frac{2\pi^{2}\Phi_{r}}{\beta}(1-\frac{4}{\gamma_{-}}(\alpha_{+}+n)^{2}-\frac{4}{\gamma_{+}}(\alpha_{-}+m)^{2})}}}{(1-4(\alpha_{+}+\alpha_{-}+m+n)^{2})(1-4(\alpha_{+}-\alpha_{-}-m+n)^{2})}.$
The power of temperature appearing in the one-loop determinant comes from the fact that $D(2,1|\upalpha)$ has $8$ fermionic and $9$ bosonic generators. The rest of the first line comes from the two $\text{SU}(2)$ degrees of freedom and the fermions. The second line includes the classical action, and a denominator that removes fermion zero-modes from the one-loop determinants. Besides the Schwarzian coupling $\Phi_{r}$ the theory depends on one continuous parameter, since $\gamma_{+}+\gamma_{-}=1$ [^13] In section 2.3, we wrote this in terms of $\upalpha$ rather than introducing separate gauge couplings $\gamma_{\pm}$., which we can take to be their ratio $\upalpha=\gamma_{-}/\gamma_{+}$.
@@
with the index $A$ and $\dot{A}$ transforming in the bifundamental of $\text{SU}(2)_{+}$ and $\text{SU}(2)_{-}$, respectively. We can pick a basis such that $Q_{1\dot{B}}$ ($Q_{2\dot{B}}$) has eigenvalue +1/2+1/2 (1/2-1/2) with respect to $J_{+}{}^{3}$, and similarly for the second index. We can also describe the $\text{SU}(2)$ generators in the same language as $J_{A}{}^{B}=(\sigma^{i})_{A}{}^{B}J_{+}^{i}$ and $J_{\dot{A}}{}^{\dot{B}}=(\sigma^{i})_{\dot{A}}{}^{\dot{B}}J_{-}^{i}$, where $\sigma^{i}$ are the Pauli matrices.
@@
$\times\sinh\left(2\pi\sqrt{2\Phi_{r}\left(E-\frac{\upalpha j_{+}^{2}+j_{-}^{2}}{2\Phi_{r}(1+\upalpha)}\right)}\right),$
@@
To describe the allowed spins of the BPS states more explicitly, consider multiplets with fixed total spin $j_{+}+j_{-}=j$ and work with one value of $j$ at a time. For fixed $j$ we have $2j+1$ ways of distributing the total spin between the two $\text{SU}(2)$ generators, namely
@@
We find that for a given $j$ there is a unique choice of $(j_{+},j_{-})$ carrying all BPS states, but the precise way the total spin is distributed among the two $\text{SU}(2)$ spins depends on the parameter $\upalpha$. For $\upalpha\sim 0$ our expression for $R_{\rm BPS}$ implies that the only multiplets with BPS states can be $(j_{+}=j,j_{-}=0)$. When we start raising $\upalpha$ there is a jump at $\upalpha=1/(2j)$ since the combination that appears in $R_{\rm BPS}$ approaches its upper bound. At $\upalpha=1/(2j)$ the representation with BPS states jumps from $(j,0)$ to (j1/2,1/2)(j-1/2,1/2). As we continue to increase $\upalpha$ we encounter a finite number of jumps at
@@
This happens until we reach $\upalpha=2j$ and we transition from (1/2,j1/2)(1/2,j-1/2) to $(0,j)$ at $\upalpha=2j$. For any $\upalpha>2j$ the BPS spin is $(0,j)$. We emphasize that for a fixed value of $\upalpha$ and fixed $j=j_{+}+j_{-}$ there is a unique BPS spin. Finally, we can check that the transition values $\upalpha_{*}$ take place precisely at the zeros of (66).